DragonFlights, Inc. is planning on purchasing a new flying dragon for their new route to Volantis. The cost of the dragon is $10.3 million. On average the dragons are operational for about 10 years. Dragon will be retired after 10 years with no salvage value. Marketing department of DragonFlights has conducted consumer surveys and market research in order to determine market demand and estimate future sales. This research has cost them $100,000. Based on the results of market studies management of the company estimates that they can make 150 flights per year on the new route the 1st year of operations. Each flight will generate sales of $55,000 per flight. Revenues are expected to decline 2% per year as the dragon gets older and doesn’t have the energy to make as many flights per year. Annual maintenance of the dragon will cost the company $150,000. Salary of the dragon pilot is $250,000 per year. Variable costs per flight are $15,000. Feeding the dragon will cost the company $450,000 per year. Other fixed costs per year are $1.25 million. Tax rate is 25%. DragonFlights has reset criteria for approving capital projects. Payback period has to be four years or less, NPV has to be positive and IRR has to beat the hurdle rate. DragonFlights uses their cost of capital as hurdle rate. DragonFlights is currently financed with common equity and debt in the form of bonds. They have 1 million shares outstanding and shares are trading at a price of $116 per share. Stock beta is 2.8 as dragon flying business is a risky business. Company is also financed with publicly traded bonds in the total amount of $50 million in book value with 50 thousand bonds outstanding. Bonds are trading at premium at for $1,250 per bond. Bonds have 20 years to maturity and have a coupon rate of 14%, with coupon paid semi-annually. Bonds have a face value of $1,000. Current risk-free rate is 2.5% and market risk premium is 7%. Marginal tax rate is 25%.

1. What proportion of each source of capital does DragonFlights use?

2. What is the cost of equity for DragonFlights?

3. What is the cost of debt for DragonFlights?

4. What is depreciation per year using straight line depreciation over 10 years and zero salvage value?
Answer #1

solution:

1:

worth of equity E =share price x no of shares =116 x 1 million =116 million

worth of debt D=no of bonds x bond price =50000 x 1250 = 62.5 million

Total worth V= 116+62.5 =178.5

Proportion of equity = E/V =116/178.5 =64.98%

Proportion of mortgage =D/V= 62.5/178.5 =35.01%

2:

price of equity = risk free rate+ beta x market risk premium

=2.5+2.8x7 =22.1%

3:

value of mortgage =YTM=10.9%

value of mortgage is the YTM of the bond which can be estimated here

Bond cashflows
Years 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40
amount 1250
Coupon payment 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70
Par worth 1000
Total cashflows -1250 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 70 1070
IRR 5.45%
YTM= 2*IRR 10.90%